# CELA: Cost-Efficient Language Model Alignment for CTR Prediction

### v2

@@ 5.6. Further Analysis
5.6.1. Compatibility for language models.
We compare the performance of CELA equipped with PLMs of various sizes: BERTTiny (14.5M) (Jiao et al., 2019), BERTSmall (29.1M) (Bhargava et al., 2021), BERTMedium (41.7M) (Turc et al., 2019), RoBERTaBase (110M) (Liu et al., 2019), BERTLarge (336M) (Devlin et al., 2018), and OPT (1.3B) (Zhang et al., 2022). For the OPT model, Causal Language Modeling (CLM) pre-training loss and LoRA fine-tuning (Hu et al., 2022) are implemented. Experimental results are illustrated in Figure 4, leading to the following insights:
  • CELA, when enhanced with a PLM, no matter its size, exhibits significant performance gains over DIN, validating that the use of PLMs can extract knowledge from textual features and benefit CTR prediction.
  • As the complexity of the comprehensive PLM escalates, the marginal gains in CELA’s performance exhibit a plateauing trend. Notably, BERTLarge underperforms relative to RoBERTaBase, attributable to the former’s significantly larger parameter set, which does not fine-tune as effectively with limited items. This observation implies that beyond a specific threshold, the addition of parameters ceases to yield proportional improvements, suggesting that a medium-sized PLM is sufficient to attain optimal performance.
  • CELA with OPT exhibits a decrease in performance compared to comprehensive PLMs. This reduction is not only because the extensive parameter space of such models proves challenging to fine-tune effectively with a limited item dataset, but it is also likely due to the generative PLM’s unsuitability as a text encoder for CTR prediction.
@@
5.6.2. Compatibility for ID-based models.
Given the orthogonal nature of our proposed framework to ID-based models, we explore its efficacy across a variety of backbones, such as AutoInt, DCNv2, DeepFM, and DIN. As shown in Table 5, CELA consistently achieves significant performance improvements in comparison to these backbones, thereby confirming the framework’s model-agnostic nature and extensive compatibility.
@@
5.6.3. Alignment of Popular Items Only
ID-based models exhibit poor performance in cold-start scenarios, suggesting inadequacies in learning item embeddings for cold-start items. We aim to ascertain whether aligning poorly learned item embeddings during the ROMA stage might lead to a decline in performance. To this end, we design an experiment in which CELA’s ROMA stage aligns only the top 20%, 40%, 60%, 80%, and 100% of items by popularity. Figure 5 indicate that CELA’s performance reaches its apex when aligning 80% of the most popular items. This suggests that limiting the alignment to a subset of popular items facilitates a more accurate mapping of item text representations to the collaborative filtering space, while aligning poorly learned embeddings will lead to performance degradation.
Figure 5. Performance with alignment of popular items only, evaluated on Amazon. The $x$-axis represents the top percentile of popular items.
@@ 5.7. Online A/B Testing
To evaluate the effect of CELA on real-world scenarios, we deploy CELA and conduct an online A/B test in an industrial Android AppGallery, which is one of most popular App stores of China.
@@ 5.7.1. Deployment
The online deployment process is outlined as:
  • Data Preprocessing. Utilizing a LLM, we refine the collected app descriptions to eliminate extraneous details, such as contact information, thereby enhancing conciseness. The refined text is indexed by the item identifier for low-cost storage.
  • Update Process. Given the infrequent changes in app descriptions, weekly training is conducted during the DAP stage. Thanks to the efficiency of the ROMA stage, it is feasible to align the PLM daily with the item embeddings from the previous day’s recommendation model. This alignment supplies the MF2 stage with the text embedding table, minimizing additional training overhead. Accordingly, updates to the MF2 stage also occur daily.
  • Online Services. The online recommendation model retrieves text representations directly from the pre-computed text embedding table via item identifiers, circumventing the need for real-time PLM text processing.
This deployment facilitates iterative and daily enhancements of both the ROMA and MF2 stages, while the efficient and effective online serving ensures optimal service delivery.
@@ 5.7.2. Online Experimental Results
DTR Impr eCPM Impr Latency
0.92% 0.97% 69.73ms
An online A/B test of the CELA is conducted, utilizing 1% of AppGallery’s extensive traffic—which includes tens of millions of daily active users—over a week. The treatment group implements CELA’s three-stage framework, extending the control group.
Table 6 reveals CELA’s significant impact: a 0.92% increase in Download-Through Rate (DTR) and a 0.97% rise in Effective Cost Per Mille (eCPM), indicating notable improvements in user engagement and revenue. Importantly, the online serving latency of CELA, complying with the scenario’s standard of under 100ms and compared to the control group’s 60.04ms, affirms its suitability for real-world applications.
@@ 6. Conclusion
In this study, we introduced a novel cost-efficient framework for CTR prediction that leverages the strengths of PLMs and textual features alongside ID-based models. To tailor the capabilities of PLMs for recommendation, we implemented domain-adaptive pre-training and item-level alignment, concurrently addressing the inefficiencies associated with PLM training at the interaction level. By abolishing hard prompts and employing PLMs solely as text encoders, we intricately describe user behaviors while preventing token overflow due to the increase in features. Our framework is model-agnostic and industry-friendly, with low training overhead and serving latency. Comprehensive experiments conducted on both public and industrial datasets validate CELA’s notable enhancements in performance.

### v3

@@ 5.6. Further Analysis
5.6.1. Compatibility for language models
We compare the performance of CELA equipped with PLMs of various sizes: BERTTiny (14.5M) (Jiao et al., 2020), BERTSmall (29.1M) (Bhargava et al., 2021), BERTMedium (41.7M) (Turc et al., 2019), RoBERTaBase (110M) (Liu et al., 2019), BERTLarge (336M) (Devlin et al., 2019), and OPT (1.3B) (Zhang et al., 2022). For the OPT model, Causal Language Modeling (CLM) pre-training loss and LoRA fine-tuning (Hu et al., 2022) are implemented. Experimental results are illustrated in Figure 4, leading to the following insights:
  • CELA, enhanced with a PLM regardless of size, significantly outperforms DIN. This validates that PLMs can effectively extract knowledge from item textual features, benefiting CTR prediction.
  • As the complexity of the comprehensive PLM increases, CELA’s performance gains plateau. Notably, BERTLarge underperforms compared to RoBERTaBase due to its larger parameter set, which does not fine-tune effectively with limited items. This suggests that beyond a certain threshold, additional parameters do not yield proportional improvements, indicating that a medium-sized PLM is sufficient for optimal performance.
  • Generative decoder-based large language models do not enhance CELA as effectively as comprehensive PLMs. Specifically, CELA with OPT shows decreased performance compared to comprehensive PLMs. This reduction is due to the difficulty in fine-tuning large parameter models with limited data and the generative PLM’s unsuitability as a text encoder for CTR prediction.
@@ 5.6.1. Compatibility for language models
5.6.2. Compatibility for ID-based models
Given the orthogonal nature of our proposed framework to ID-based models, we explore its efficacy and efficiency across a variety of backbones. As shown in Table 4 and Appendix C.1, CELA consistently achieves significant performance improvements and comparable inference and training costs when compared to these backbones, thereby confirming the framework’s model-agnostic nature and extensive compatibility.
@@ 5.6.2. Compatibility for ID-based models
5.6.3. Analysis of Alignment w.r.t. Popular Items
Figure 5. Performance with alignment w.r.t. popular items, evaluated on Amazon. The $x$-axis represents the top percentile of popular items.
ID-based models exhibit poor performance in cold-start scenarios, suggesting inadequacies in learning ID embeddings for cold-start items. We aim to ascertain whether aligning poorly learned ID embeddings during the ROMA stage might lead to a decline in performance. To this end, we design an experiment in which CELA’s ROMA stage aligns only the top 20%, 40%, 60%, 80%, and 100% of items by popularity. Figure 5 indicates that CELA’s performance reaches its apex when aligning 80% of the most popular items. This suggests that limiting the alignment to a subset of popular items facilitates a more accurate mapping of item text representations to the collaborative filtering space while aligning poorly learned ID embeddings will lead to performance degradation.
@@ 5.7. Online A/B Testing
To evaluate the effect of CELA on real-world scenarios, we deploy CELA and conduct an online A/B test in an industrial app advertising platform, which is one of the most popular app stores.
@@ 5.7.1. Deployment
The industrial app advertising platform, illustrated in Figure 6, comprises two main components: offline training and online serving. In offline training, an LLM refines the collected app textual features from the app pool, removing extraneous details such as contact information to enhance conciseness. The clean text is indexed by the app identifier for low-cost storage. Given the infrequent changes in app textual features, training is conducted weekly during the DAP stage. The efficient ROMA stage enables daily alignment of the PLM with ID embeddings from the current ID-based backbone, providing the MF2 stage with the text embedding table and minimizing additional training overhead. Consequently, updates to the MF2 stage also occur daily. In online serving, when a user initiates a request, the candidate apps are scored by CELA, which retrieves the text embeddings of candidate apps directly from the pre-computed text embedding table, bypassing real-time text processing. The candidate apps are then ranked according to Effective Cost Per Mille (eCPM) and the predicted Download-Through Rate (DTR) scores, with the top-$k$ apps displayed in the list.
Figure 6. Overview of the app advertising platform.
@@ 5.7.2. Online Experimental Results
eCPM DTR Training Time Inference Time
+1.48% +0.93% -19.41% +9.69ms
An online A/B test of CELA is conducted on the advertising platform over three weeks, including tens of millions of daily active users. The treatment group implements CELA’s three-stage framework, extending the control group. Each group is trained in a single cluster, where each node contains an 8-core Intel(R) Xeon(R) Gold 6151 CPU (3.00GHz), 32GB RAM, as well as 1 NVIDIA TESLA V100 SXM2 GPU with 32GB memory.
Table 5 shows the superiorities of CELA: a 1.48% rise in eCPM and a 0.93% increase in DTR, indicating a positive influence on user engagement and revenue. Notably, the latency of CELA is below the 100ms threshold , which is acceptable for real-world applications.
Figure 7. Performance comparison w.r.t. unpopular apps. The bar graph represents the exposure for each group, while the line chart represents the relative improvement compared to the control group.
To explore the impact of CELA on cold-start items in the real business scenario, we group apps according to baseline exposure and measure the relative improvements in eCPM and DTR. Each group consists of an equal number of apps, focusing on the 15 least popular groups as illustrated in Figure 7. The results demonstrate that CELA significantly improves performance in unpopular groups. This enhancement is credited to the PLM that leverages external knowledge to create more accurate item profiles.
@@ 6. Conclusion
In this study, we introduced a novel cost-efficient framework for CTR prediction that leverages the strengths of PLMs and textual features alongside ID-based models. To tailor capabilities of PLMs for recommendation, we implemented domain-adaptive pre-training and item-level alignment, concurrently addressing the inefficiencies associated with PLM training at the interaction level. By abolishing hard prompts and employing PLMs solely as text encoders, we intricately describe user behaviors while preventing token overflow. Our framework is model-agnostic and industry-friendly, with low training overhead and serving latency. Comprehensive experiments conducted on both public and industrial datasets validate CELA’s notable enhancements in performance.
Appendix A Corpus Construction
For illustration, consider the features of an item detailed as follows:
    "item_id": 31852,
    "brand": "Coxlures",
    "categories": ["Sports & Outdoors", "Dance"],
    "description": "The ballet skirt is nice for
                    dress-up play for ballerinas."
Categorical features such as item_id, brand, and categories can be encoded using conventional methods like one-hot or multi-hot encoding. However, textual features like the description, require language models for encoding. To enhance the language model’s understanding of dataset-specific texts, a corpus is constructed from item descriptions. This corpus is organized as follows:
    item 1’s description,
    item 2’s description,
    item |I|’s description
Appendix B Experimental Settings
B.1. Datasets
Offline experiments are conducted on the datasets below. The statistics are presented in Table 6.
Table 6. Statistics of datasets for offline experiments.
Dataset #Users #Items #Features #Samples
Amazon 300,562 100,682 6 4,494,230
MovieLens 92,911 26,084 4 10,146,329
AppGallery 11,056,819 29,759 8 16,398,193
[…]
MovieLens Dataset  [^2] https://grouplens.org/datasets/movielens/latest/ collects the rating activity from MovieLens  [^3] http://movielens.org/. We extract movie summaries as the textual feature from provided hyperlinks and sample 10 million interactions through random sampling of users. The task is to predict whether a user will like the next item, indicated by a rating of 3 or higher, based on a sequence of up to 20 user behaviors.
AppGallery Dataset comprises application exposure and download records sampled from an industrial app advertising platform over eight days with user consent. Data from the first seven days constitute the training set, while data from the eighth day serve as the test set. The task is to predict whether the user will download the next app, using app descriptions as the textual feature and a sequence of up to 14 user behaviors.
B.2. Implementation Details
In our offline experiments, models are implemented using RecStudio (Lian et al., 2023) and Transformers (Wolf et al., 2020). We integrate RoBERTa (Liu et al., 2019) to enhance the optimal ID-based model. AdamW (Loshchilov and Hutter, 2017) optimizer facilitates all training stages. For the DAP stage, $\tau_{1}$ ranges from [0.001, 0.01, 0.05, 0.1] and $\alpha$ from [0.1, 0.5, 1.0, 1.5], with a batch size of 32. In the ROMA stage, $\tau_{2}$ is 1, batch size is 128, the language model’s learning rate is 5e-5, and projection units are [128, $f_{i}d$], with $f_{i}$ as item-side feature count and $d$ set at 16. The MF2 stage uses a batch size of 1024 and standardized hidden units [128, 64, 64] across ID-based models. ID-based models that do not focus on user behavior modeling aggregate user behavior sequences using mean pooling.
Appendix C Experimental Results
C.1. Compatibility for ID-based models
We extend the evaluation of CELA’s model-agnostic nature by testing it with a wider range of ID-based models and providing a detailed analysis of the computational costs associated with each integration. The results are presented in Table 7. It is worth noting that CELA’s training involves both the PLM and the backbone, with the PLM requiring 1.07 hours for pre-training and weekly updates, while the backbone is updated daily.
Table 7. Training and inference efficiency on AppGallery. The inference time is calculated for one batch and the training time is for the whole process.
Model Inference (ms) Training (h)
AutoInt 62.83 2.90
$\text{CELA}_{\text{AutoInt}}$ 77.96 6.99
DCNv2 59.63 2.28
$\text{CELA}_{\text{DCNv2}}$ 64.16 5.66
DeepFM 62.67 2.88
$\text{CELA}_{\text{DeepFM}}$ 65.36 4.77
xDeepFM 60.96 2.33
$\text{CELA}_{\text{xDeepFM}}$ 65.36 4.85
DIN 61.38 1.76
$\text{CELA}_{\text{DIN}}$ 69.64 5.42
DIEN 66.94 2.93
$\text{CELA}_{\text{DIEN}}$ 75.78 5.86